wiki:PT1_2Report20110515
Last modified 8 years ago Last modified on 06/01/2011 11:54:14 AM

Table of Contents

    Previous month PT1.2 - Mgt/SE

    Current month Mgt/SE

    Next month PT1.2 - Mgt/SE

    Guide on Reporting Percentage Completed:

    For software development tasks:

    • use 0% if the solution is still being designed;
    • use 25% if the solution's design is completed;
    • use 50% if the solution is implemented and tested within the Ticket branch;
    • use 75% if the implementation review is complete, the Ticket branch has been merged onto the Trunk branch, and the Ticket is closed.
    • use 100% if the affected Trunk packages are tagged and released.

    For level of effort tasks, use % of effort expended rounded to nearest 5%:

    For other tasks, still use these increments, but apply their meaning as best as possible:

    • use 0% if the deliverable is still being designed;
    • use 25% if the deliverable's design is completed;
    • use 50% if the deliverable is implemented awaiting review/approval;
    • use 75% if the deliverable has been reviewed and in revision;
    • use 100% if the deliverable is approved and posted to appropriate repository.

    Note: Please place an asterisk (*) after the percent if you updated this value in this report. Please update the "Last Update" date even if the value doesn't change, so we'll know whether we have the most current information.

    Task Start Finish Resources (lead first) % Complete Comments Last Updated
    Infrastructure Mike Freemon
    Acquire and configure new storage to replace SAN 20110131 20110503 Freemon Mike[5%] 75% This machine has arrived, and is being integrated into our cluster. 20110517
    Relocate servers from ACB to NCSA building 20110425 20110627 Gehrig David[10%] 25% Servers will be relocated next week 20110519
    Environment and Tools Ray Plante
    TCT approval for environment/tool updates 20110309 20110418 Allsman Robyn[50%] 50% This is an on-demand task. 'eigen 3.0 and SOFA' experimental use was OKed. 20110516
    Install, configure, publish environment/tool updates 20110201 20110422 Plante Ray[25%] 25% 20110424
    Complete eups release 20110208 20110503 Plante Ray[25%] 25% 20110424
    Update environment and tools documentation 20110503 20110506 Lim KT[50%] 0% Still not sure what this is. 20110520
    Support trac, svn, systems administration 20110131 20110617 Gehrig David[15%] 70%* ongoing item 20110519
    ImSim Data Simon Krughoff
    Prepare and code to generate images and catalogs 20110131 20110505 Krughoff Simon[30%],Lim KT[5%] *75% The last part of this task is to deliver new calibration products. This should be done by next week. 20110513
    Test generation of images and catalogs 20110328 20110506 Krughoff Simon[30%] 75% Stars + Galaxies Tuesday run data delivered 20110513
    Generate images and catalogs 20110506 20110610 Krughoff Simon[30%],Gibson Rob *50% Input catalogs are generated. First run of production data is underway. Reference catalogs are being generated 20110513
    Support ImSim data 20110610 20110708 Krughoff Simon[10%] 0% 20110513
    Import reference catalogs into Gator for comparison to output catalogs 20110610 20110614 Monkewitz Serge[50%] 50% I've been loading reference catalogs to lsst10 as I get them, Gator ingest doesn't make sense until we have reference catalogs we know we'll use for the 2000x runs 20110601
    Data Access Middleware and Database KT Lim
    Serialization and Persistence additions/fixes 20110214 20110427 Lim KT[50%] 100%* Fixes for Source "blob" persistence are not required for PT1.2. Major butler changes are complete. 20110520
    Schema updates for galaxy photometry and others as needed 20110309 20110420 Becla Jacek[25%] 100%* 20110516
    Spatial support (deploying udfs, indexes etc) 20110325 20110509 Serge Monkewitz [20%] Becla Jacek[20%] 75%* Deployed spatial UDFs on lsst10 (basic version). Serge implemented HTM-based solution (faster, more flexible, better APIs than the basic version) - doing final tests now, will deploy in 1-2 weeks. 20110516
    Improve schema documentation (udfs, stored functions/procedures) 20110509 20110606 Becla Jacek[25%] 10%* Started designing the format of documentation 20110516
    ISR and CCD Assembly Simon Krughoff
    Implement more SDQA metrics for ISR 20110418 20110429 Krughoff Simon[50%] 0% 20110513
    Clean-up/re-write ISR python code 20110323 20110517 Krughoff Simon[50%] 25% Begun as part of the #1556 merge 20110513
    Implement raw CCDAssembly method 20110517 20110524 Krughoff Simon[50%] 0% 20110513
    Point Source Photometry Robert Lupton
    Identify and fix zeropoint calculation error 20110131 20110418 Becker Andy[25%],Lupton Robert[25%] 100% This appears solved, modulo addressing vignetting at the edges of the focal plane 20110516
    Test aperture corrections 20110204 20110420 Lupton Robert[25%] 75% 20110421
    PSF Estimation Robert Lupton
    Finish PSF unification 20110131 20110427 Owen Russell[20%],Jarvis Mike[80%] 25% Merged HSC changes into meas_algorithms ticket #1573 and updated meas_pipeline to match. Mike's code still needs unit tests. 20110516
    Update PSF formatter 20110418 20110425 Lim KT[50%] 0% Depends on previous 20110520
    Debug PCA PSF code 20110131 20110503 Lupton Robert[50%] 75% 20110420
    Galaxy Photometry Jim Bosch Code is working in small-scale tests. Still need to work on integration into the source measurement stage code. 20110516
    R&D for optimizers, parameterizations, and inputs 20110415 20110506 Dubcovsky Martin[50%],Bosch Jim[50%] 75% Implmentation complete. Still need to do integration testing and tune algorithmic options. 20110516
    Implement Kron photometry 20110214 20110419 Lupton Robert[25%] 75% Not critical for production start 20110419
    Fix Source class for memory bloat for Source Association 20110524 20110531 Monkewitz Serge[25%],Lupton Robert[50%] 0% Fixing this properly is a post 1.2 item; initial galaxy photometry runs will be used to determine if any immediate action is needed 20110601
    Implement Chebyshev Function 20110419 20110421 Owen Russell[50%] 100%* 20110516
    Update SourceAssoc to average galaxy photometry measurements 20110601 20110630 Monkewitz Serge 0% For now, this will be limited to fluxes only, shapes may be included at a later date 20110601
    Co-addition Owen Russell 20110516
    Pipette scripts 20110131 20110420 Owen Russell[50%] 50% All scripts are written and updated for the #1556 merge but those changes need to be tested and outlier-rejected coadd is new and completely untested. 20110516
    Write outlier-rejected psf-matched co-addition 20110315 20110502 Owen Russell[50%] 25% Written as a pipette script but not tested. No pipeline stage in coadd_pipeline yet. 20110516
    Chi-squared co-addition 20110329 20110503 Owen Russell[50%] 75% coadd_chisquared package is done. Needs a pipette script and coadd_pipeline stage updates. 20110516
    Update co-addition to use the new afw with unified geometry classes 20110404 20110504 Owen Russell[50%] 75% coadd_utils and coadd_chisquared are ready. coadd_psfmatched has been udpated for #1556 but not tested (I had to update ip_diffim first, now I'm busy with other things). coadd_pipeline needs updating (but is not used by pipette coaddition scripts). 20110516
    Finish sky pixelization 20110504 20110601 Owen Russell[50%] 0% 20110516
    Templates and Difference Imaging Becker Andy 20110516
    Merge ip_diffim development ticket to trunk 20110131 20110506 Becker Andy[50%] 50% Waiting on #1649, #1668. Merging in changes from #1648. 20110516
    Support Psf matching for template creation 20110228 20110513 Becker Andy[50%] 50% 20110516
    Assess quality of generated templates 20110516 20110610 Becker Andy[50%] 0% No templates generated yet 20110516
    Implement image subtraction on visit pairs 20110330 20110513 Becker Andy[50%] 50% Requires merging #1140 to trunk 20110516
    Quality Assessment Andy Becker 20110516
    Implement SDQA metrics for Source Association 20110309 20110418 Monkewitz Serge 90% Implemented reference catalog to source match as part of SourceAssoc ingest 20110601
    Implement QA Pipeline package 20110131 20110513 Bickerton Steve[50%],Becker Andy[5%] 50% Is being run on all weekly runs. Bickerton is currently working on documentation. 20110516
    Generate automatic QA metrics on pipeline runs 20110328 20110530 Becker Andy[50%] 75% Metrics are being generated, tho its unclear if anyone is using them yet. 20110516
    Build and Test Robyn Allsman 75% Buildbot is operational - automatically testing Release and Development versions on a nightly basis and allowing on-demand testing via a web-form interface. 20110516
    Install and integrate new Parasoft release 20110131 20110420 Allsman Robyn[50%] 50% Parasoft is debugging their recent release which no longer handles LSST's C++ template usage; expected in June/July? timeframe 20110516
    Buildbot development for pipette 20110420 20110427 Allsman Robyn[50%] 0% No action 20110416
    Incorporate standards into buildbot trunk validation 20110427 20110713 Allsman Robyn[25%] 0% No action 20110516
    Maintain buildbot integration and test 20110131 20110722 Allsman Robyn[25%] 50%* See daily status page: http://dev.lsstcorp.org/trac/wiki/DM/buildbot/Daily_Status 20110516
    Production Runs Greg Daues
    Migrate processing from abe to new platform 20110225 20110513 Plante Ray[10%],Daues Greg[10%] 50% Production scale runs (189 workers - 16 nodes on a limited data set [506 CCDs], 378 workers - 32 nodes on 506 CCDs) have executed without apparent errors on TACC Ranger. Small scale runs are also succeeding on TACC Lonestar. On both platforms the issue of getting a consistent stack of up-to-date tagged releases is ongoing. 20110519
    Weekly runs 20110323 20110511 Allsman Robyn[5%] 60%* Maintaining/updating the weekly runs is a continuing process, e.g., changes to the run characteristics such as: new input data, new reference catalogs, and new hardware. 20110516
    Production debug/practice runs 20110420 20110511 Daues Greg[20%] 0% 20110421
    Full Production runs 20110511 20110629 Daues Greg[50%] 0% 20110421
    User Support Dick Shaw
    Update Data Challenge Handbook 20110131 20110624 Shaw Dick[50%] 25% This will not be done before Production. 20110518
    Update DM User Support 20110421 20110617 Shaw Dick[50%] 25% Includes SIA tools evaluation, setup SIAP, REDDnet image cache testing 20110518
    Configure/tune catalogs 20110526 20110624 Becla Jacek[10%] 50% All tune-ups needed so far has been done 20110421
    Data Analysis Tim Axelrod
    DM Scientist/Engineer? data analysis 20110420 20110629 Axelrod Tim[50%],Lupton Robert[50%],Bosch Jim[50%],Becker Andy[50%],Van Dyk Schuyler[25%],Shaw Dick[50%] 0% 20110421
    Science Collaboration data analysis 20110627 20110805 Science Collaboration 0% 20110421
    Data analysis report in handbook 20110808 20110819 Axelrod Tim[50%],Lupton Robert[25%],Bosch Jim[50%],Becker Andy[50%],Van Dyk Schuyler[25%],Shaw Dick[50%],Science Collaboration 0% 20110421